OpenStudy (anonymous):

2=1/x+4+2 I got x = -1/4 Am I right???

OpenStudy (cj49):

wats the question

jhonyy9 (jhonyy9):

i think x=?

OpenStudy (anonymous):

I was asking if I was right when I solved 2=1/x+4+2. I got x = -1/4. Was I correct or not?

OpenStudy (anonymous):

Is this your question? \[2 = \frac{1}{x+4} + 2\]

jhonyy9 (jhonyy9):

so there are 1 -- + 4 +2 ? x

OpenStudy (anonymous):

@waterineyes, yes.

OpenStudy (anonymous):

Then 2 and 2 will simply get cancelled.. You are left with: \[0 = \frac{1}{x+4}\]

OpenStudy (anonymous):

So I was correct?

OpenStudy (anonymous):

I never said that..

OpenStudy (anonymous):

Okayyy...

OpenStudy (anonymous):

Where we have reached?? This way you will get as Undefined..

OpenStudy (anonymous):

Look for your question once again..

OpenStudy (anonymous):

You must be kidding.. Your question should be: \[2 = \frac{1}{x} + 4 + 2\]

OpenStudy (anonymous):

Yeah, that was it. I screwed up. After all, nobody is perfect, eh?

OpenStudy (anonymous):

@jhonyy9 was right then..

OpenStudy (anonymous):

Subtract 2 from both the sides first.. then subtract 4 from both the sides, and then YES, you will get what you got.. :P

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

Woo hoo :)